# Theory of Mind in Large Language Models: Assessment and Enhancement

### v1

@@ Appendix D Comparative Information on Benchmarks
FANTOM LLM-generated conversations related to pets, personal growth, and travelling etc Conversation Open-Ended, Yes/No, Multi-Choice QA 2 English EMNLP (2023)
@@ Appendix E Interactive Benchmarks
Although they are not yet widely used for evaluating LLMs’ ToM capabilities, a few interactive benchmarks or tasks have been proposed to help improve or test neural models’ ToM skills. Most of these benchmarks are derived from game-based or reinforcement learning research scenarios, such as Minecraft Bara et al. (2021), Dungeons and Dragons Zhou et al. (2023b), and Grid Rabinowitz et al. (2018); Sclar et al. (2022); Ma et al. (2023b). These benchmarks can be adapted to develop situated and interactive evaluations for LLMs. A few potential interactive benchmark candidates are listed below.
@@
Situated ToM in Grid World Ma et al. (2023b). Ma et al. (2023b) advocate for a situated evaluation of Theory of Mind (ToM), as it encompasses more aspects than text-only datasets while also mitigating data contamination and shortcuts. They compile a multiple-choice question-answering dataset in MiniGrid Chevalier-Boisvert et al. (2018), which covers all aspects of ATOMS as well as a "reality check."
Beyond these benchmarks, there is still room for further improvement. For example, incorporating more agents, as the current MINDCRAFT Bara et al. (2021) and G-DRAGON Zhou et al. (2023b) benchmarks involve only two; integrating multimodal information, as G-DRAGON Zhou et al. (2023b), and Situated ToM in Grid World Ma et al. (2023b) currently focus solely on text input; and extending introspective beliefs to first-order or even higher-order beliefs Li et al. (2023). Additionally, negotiation-based or tabletop games, such as auction games, and human-AI collaboration tasks Zhang et al. (2024a), could further contribute to benchmark development for evaluating the ToM capabilities of LLMs.
Approaches Order Fine-Tuning
SYMBOLICTOM Sclar et al. (2023) Designed for any depth, tested up to third-order No
SIMTOM Wilf et al. (2024) Tested up to second-order No
PercepToM Jung et al. (2024) Tested up to second-order No
TIMETOM Hou et al. (2024) Designed for any depth, tested up to third-order No
ToM-LM Tang and Belle (2024) Tested up to second-order Yes
BIP-ALM Jin et al. (2024) Tested only first-order Yes
LIMP Shi et al. (2024) Tested up to third-order No
@@ Appendix G Additional Future Directions
Exploring multilingual benchmarks. Currently, most of the benchmarks are still in English. Multilingual benchmarks Chen et al. (2024) are crucial for evaluating and advancing LLMs’ comprehensive Theory of Mind capabilities across different languages.
Advancing higher-order belief reasoning: Most existing strategies and benchmarks still focus on second-order beliefs, whereas higher-order belief reasoning demands further attention and innovation Wu et al. (2023); Hou et al. (2024) in order to achieve "advanced ToM" Białecka-Pikul et al. (2017); Sclar et al. (2023) in LLMs.
Exploring the ToM capabilities in smaller language models: Most strategies have been tested on models with over 7 billion parameters. However, these strategies may not be effective on models with fewer than 7 billion parameters, as smaller language models typically have a reduced capacity for understanding instructions Hou et al. (2024). Additionally, LLMs require substantial resources, which may not always be available. Therefore, it is crucial to investigate the ToM capabilities of smaller language models. Appendix F presents several effective methods developed during the era of small language models that may offer valuable insights for this research direction.

### v2

@@ Appendix D Comparative Information on Benchmarks
FANTOM LLM-generated conversations related to pets, personal growth, and travelling etc. Conversation Open-Ended, Yes/No, Multi-Choice QA 2 English EMNLP (2023)
@@ Appendix E Interactive Benchmarks
Although they are not yet widely used for evaluating LLMs’ ToM capabilities, a few interactive benchmarks or tasks have been proposed to help improve or test neural models’ ToM skills. Most of these benchmarks are derived from game-based or reinforcement learning (RL) research scenarios, such as Minecraft Bara et al. (2021), Dungeons and Dragons Zhou et al. (2023b), and Grid Rabinowitz et al. (2018); Sclar et al. (2022); Ma et al. (2023b). These benchmarks can be adapted to develop situated and interactive evaluations for LLMs. A few potential interactive benchmark candidates are listed below.
@@
Situated ToM in Grid World Ma et al. (2023b). Ma et al. (2023b) advocate for a situated evaluation of ToM, as it encompasses more aspects than text-only datasets while also mitigating data contamination and shortcuts. They compile a multiple-choice question-answering dataset in MiniGrid Chevalier-Boisvert et al. (2018), which covers all aspects of ATOMS as well as a "reality check."
Beyond these benchmarks, there is still room for further improvement. For example, incorporating more agents, as the current MINDCRAFT Bara et al. (2021) and G-DRAGON Zhou et al. (2023b) benchmarks involve only two; integrating multimodal information, as G-DRAGON Zhou et al. (2023b), and Situated ToM in Grid World Ma et al. (2023b) currently focus solely on text input; and extending the method from handling introspective beliefs to addressing first-order or even higher-order beliefs questions Li et al. (2023). Additionally, negotiation-based or tabletop games, such as auction games, and human-AI collaboration tasks Zhang et al. (2024a), could further contribute to benchmark development for evaluating the ToM capabilities of LLMs.
@@ Appendix G Additional Future Directions
Exploring multilingual benchmarks. Currently, most of the benchmarks are still in English. Multilingual benchmarks Chen et al. (2024) are crucial for evaluating and advancing LLMs’ comprehensive ToM capabilities across different languages.
Advancing higher-order ToM reasoning. While most existing strategies and benchmarks remain focused on second-order beliefs, higher-order reasoning is crucial in contexts such as multi-party negotiations. As such, advancing higher-order ToM capabilities warrants greater attention and innovation Wu et al. (2023); Hou et al. (2024) to achieve "advanced ToM" Białecka-Pikul et al. (2017); Sclar et al. (2023) in LLMs.
Exploring the ToM capabilities in smaller language models. Most strategies have been tested on models with over 7 billion parameters. However, these strategies may not be effective on models with fewer than 7 billion parameters, as smaller language models typically have a reduced capacity for understanding instructions Hou et al. (2024). Additionally, LLMs require substantial resources, which may not always be available. Therefore, it is crucial to investigate the ToM capabilities of smaller language models. Appendix F presents several effective methods developed during the era of small language models that may offer valuable insights for this research direction.